# Enhanced superconductivity in LaRu3Si2 by chemical pressure tuning of kagome flat bands

### v1

@@ Abstract
In kagome metals, flat bands induced by frustrated hopping serve as a platform for strong electronic correlation, while often exhibiting substantial dispersion along the $k_{z}$ direction due to interlayer coupling. Here, we investigate the superconductivity of kagome metal LaRu3(Si1-xGex)2 by chemical pressure tuning while preserving the Ru-$4d$ states that constitute the kagome flat bands. We observe a sizable enhancement in the density of states up to $x=0.07$, as determined by the specific heat, with a concomitant increase in the superconducting transition temperature. As supported by our first-principles calculations, the lattice expansion along the $c$-axis by chemical pressure mitigates a detrimental effect of hybridization between kagome layers, thereby reinforcing the three-dimensional localization of flat-band electrons. Providing evidence for LaRu3Si2 as a flat-band hosting superconductor, we demonstrate a simple, yet effective, route to engineer flat bands by uniaxial lattice expansion.
Main Text
@@
The kagome lattice, a two-dimensional network of corner-sharing triangles, features three sublattice degrees of freedom per unit cell. kagome metals serve as a platform for investigating the intricate interplay between electron-electron correlations and the topology of electronic band structures. Yin et al. (2022); Wang et al. (2023); Checkelsky et al. (2024). This study discusses superconductivity in the intermetallic LaRu3Si2, where the kagome lattice is formed by Ru atoms as depicted in Fig. 1(a). Considering first a single orbital model on an isolated kagome lattice in Fig. 1(b), one obtains three bands forming a flat band and Dirac cone, and with van-Hove singularities at the $M$-points in the Brillouin zone [Fig. 1(c)]. In particular, the dispersionless upper band originates from frustrated hopping between the three sites of the kagome lattice Bergman et al. (2008). Further introducing Kane-Mele type spin-orbit coupling Kane and Mele (2005), a gap opens between the Dirac cone and the flat band, imparting a Chern number onto the flat band in this two-dimensional model Guo and Franz (2009a); Bolens and Nagaosa (2019); Kang et al. (2020).
@@
Real-world kagome metals retain the features of this simplified electronic structure if the coupling between consecutive kagome layers is weak, and mostly $3d$ electron systems have been considered as realizations of this scenario Lin et al. (2018); Kang et al. (2019, 2020); Li et al. (2021); Regnault et al. (2022); Ye et al. (2024). As an example, Fig. 1(d) presents the electronic band structure of LaRu3Si2, which hosts an hourglass-shaped dispersion originating from Ru-$4d$ orbitals along the $\Gamma$$K$$M$ path. Early experimental work and a recent theoretical study emphasize that LaRu3Si2 distinguishes itself as the only known kagome superconductor hosting flat bands near the Fermi level without magnetic order Mielke et al. (2021); Junze et al. (2025). It is further suggested that the flat bands may be a driver of superconductivity with the highest observed transition temperature $T_{\mathrm{c}}=6.7\,$K among the family of $RT_{3}M_{2}$ ($R$ = rare-earth, $T$ = transition metal and $M$ = main group element) Barz (1980); Ku et al. (1980); Chaudhary et al. (2023); Gui and Cava (2022); Rauchschwalbe et al. (1984); Mielke et al. (2021); Gong et al. (2022) and among Kagome metals as a material class Deng et al. (2025). Moreover, charge-density-wave orders, possibly accompanied by breaking of time-reversal symmetry, have been recently observed Plokhikh et al. (2024); Mielke et al. (2024), reminiscent of the kagome superconductors $A$V3Sb5 ($A=$K, Cs, and Rb) Wilson and Ortiz (2024); Ortiz et al. (2019).
@@
Real kagome metals are three-dimensional, and their flat bands often exhibit strong dispersion along the $k_{z}$ direction of reciprocal space Kang et al. (2020); Huang et al. (2022); Ye et al. (2024). This behavior is also observed in LaRu3Si2 along the $\Gamma-A$ line in Fig. 1(d). Therefore, flat-band electrons are not fully localized in known bulk kagome metals. To achieve tailored flat bands, it has thus been proposed to either suppress interlayer coupling [$t_{\mathrm{out}}$ in Fig. 1 (b)] by isolating kagome layers Jovanovic and Schoop (2022), or to introduce 3D frustrated hopping Wakefield et al. (2023). For the latter approach, 3D flat bands have been predicted for the pyrochlore lattice Guo and Franz (2009b), a 3D analogue of the kagome lattice, and have recently been observed Wakefield et al. (2023); Huang et al. (2024). Despite its simplicity, the former approach has not yet been conclusively implemented to enhance electron correlation phenomena and superconductivity on the kagome lattice.
@@
Here, we engineer the electronic structure of LaRu3Si2 by substituting Si with the larger volume element Ge. We observe a marked increase of the superconducting critical temperature $T_{\mathrm{c}}$. An enhancement of the total density of states (DOS) with Ge-doping underpins the observed evolution of $T_{\mathrm{c}}$, as confirmed by specific heat measurements. Our electronic structure calculations demonstrate that anisotropic lattice expansion along the $c$-axis isolates the kagome layers, weakens the interlayer hopping $t_{\mathrm{out}}$, and promotes the DOS at the Fermi level while keeping the intralayer hopping $t$ intact. As summarized in Fig. 1(e), this is in stark contrast to previous attempts to tune superconductivity in LaRu3Si2 by chemical substitution at the Ru site, which generally results in a sharp decrease of $T_{\mathrm{c}}$ Li et al. (2012, 2016); Chakrabortty et al. (2023).
@@
Crystals were obtained by the melting technique SI . In Fig. 2(a), the LaRu3Si2 structure was confirmed by powder X-ray diffraction (XRD). The absence of the LaRu2Si2 phase is confirmed. Figure 2(c,d) shows the hexagonal lattice constants of LaRu3(Si1-xGex)2(LRSG) extracted from XRD and their evolution with $x$ based on the LeBail analysis in the $P6/mmm$ structure. This neglects minor effects of symmetry-lowering charge order in this family of compounds Plokhikh et al. (2024). We find a robust value of $a$ and a linear expansion of the out-of-plane lattice constant $c$, which in combination gives a linear volume expansion in Fig. 2(e). The latter is consistent with Vegard’s law, and in good agreement with the trend derived from structure relaxation calculations of LaRu3Si2 and LaRu3Ge2. Magnetization and specific heat measurements were performed in a commercial cryostat SI .
@@
To confirm the bulk superconducting state in LRSG, we show the magnetic susceptibility $\chi$ in Fig. 3(a). Above the transition temperature $T_{\mathrm{c}}$, $\chi$ is small and positive, consistent with the weak paramagnetic Pauli susceptibility of LaRu3Si2 Kishimoto et al. (2002). Strong diamagnetism accompanying the superconducting transition appears as a sharp drop in the data around $6\sim 7\,$K.
@@
To precisely determine $T_{\mathrm{c}}$, we measure the specific heat $C(T)$ in zero field and at 555\,5T, which is high enough to entirely suppress the superconductivity [Fig. 3(b)]. In the former traces, a sharp kink appears at $T_{\mathrm{c}}$, the second-order transition into the superconducting state. We model the normal state contribution $C_{\mathrm{N}}$ by electronic and phonon terms, viz. $C_{\mathrm{N}}(T)=\gamma T+\beta T^{3}+\eta T^{5}$ [lines in Fig. 3(b)]. The higher order phonon contribution $\eta T^{5}$ is necessary to describe the data  Li et al. (2011). After subtraction of $C_{\mathrm{N}}(T)$, we fit the superconducting specific heat $C(T)-C_{\mathrm{N}}(T)$ to the prediction of Bardeen-Cooper-Schrieffer (BCS) theory by equal-area construction in Fig. 3(c) Tinkham (2004), confirming an enhancement in $T_{\mathrm{c}}$ with $x$. Next, we discuss the evolution of $\gamma$ and $\beta$ from the fit of $C_{\mathrm{N}}(T)$ in Fig. 3(d). A clear increase of $\gamma$ and $\beta$ is observed with $x$. Given the formula $\gamma=\pi^{2}k_{\mathrm{B}}^{2}N(E_{\mathrm{F}})/3$ with $N(E_{\mathrm{F}})$ being the density of states at the Fermi energy $E_{\mathrm{F}}$, the enhancement of $\gamma$ corresponds to an increase of $N(E_{\mathrm{F}})$ [Fig. 3(e)]. The trend is reproduced by DFT, where $N(E_{\mathrm{F}})$ increases from $2.9\,\mathrm{states/eV/u.c/spin}$ to $3.3\,\mathrm{states/eV/u.c/spin}$ as seen in Fig. 3(f). The marked enhancement of $N(E_{\mathrm{F}})$ goes hand in hand with an increase of $T_{\mathrm{c}}$, reaching up to 7.17.17.1\,7.1K for $x=0.07$ [Fig. 3(e)].
The increase of $N(E_{\mathrm{F}})$ and $T_{\mathrm{c}}$ qualitatively follows the McMillan equation
$T_{\mathrm{c}}=\frac{\Theta_{\mathrm{D}}}{1.45}\exp\left(\frac{-1.04(1+\lambda)}{\lambda-\mu^{*}(1+0.62\lambda)}\right),$ (1)
where $\Theta_{\mathrm{D}}$ is the Debye temperature, $\lambda$ is the electron-phonon coupling constant, and $\mu^{*}$ is the Coulomb pseudopotential, which usually takes a value around $0.1$ for a metal Carbotte (1990). Assuming a relatively static pairing potential $V$, the enhancement of $N(E_{\mathrm{F}})$ directly corresponds to that of $\lambda$, as given by the relation $N(E_{\mathrm{F}})V=(\lambda-\mu^{*})/(1+\lambda)$, which in turn promotes $T_{\mathrm{c}}$ SI .
The observed enhancement of $\gamma$ can be attributed to the expansion of the lattice constant $c$, as illustrated in Fig. 3(e). This interpretation is further supported by our band structure calculations for LaRu3Ge2. The projected band structures for LaRu3Ge2 and LaRu3Si2 are shown in Fig. 4(a),(b). Focusing on the $\Gamma-A$ line, which is parallel to the $k_{z}$-axis, we find that substitution of Si with Ge decreases the bandwidth of flat bands at $E_{\mathrm{F}}$. This indicates that the uniaxial lattice expansion along the $c$-axis by Ge doping suppresses interlayer coupling between kagome layers. Furthermore, the partial density of states in Fig. 4(c),(d) reveals a preference for the planar $4d_{xy}$ orbital over $4d_{z^{2}}$, with the latter exhibiting stronger hybridization along the $z$-direction. We note that the partial density of states (PDOS) for Ru-$4d_{xz}$ around $E_{\mathrm{F}}$ is hardly affected by Ge substitution SI . Importantly, Ge substitution leaves the crucial kagome lattice plane of Ru atoms mostly undisturbed, providing a powerful route for tuning superconductivity in LaRu3Si2. We stress the general applicability of such careful alloying studies to binary and ternary kagome metals Chen et al. (2024), with wide tunability of electronic properties, charge ordering, and superconductivity.
@@
Acknowledgements We thank T.-h. Arima for enlightening discussions. This work was supported by JSPS KAKENHI Grants Nos. JP21K13877, JP22H04463, JP22K20348, JP23H05431, 23K13057, 24H01607, and 24H01604, as well as JST CREST Grant Nos. JPMJCR1874 and JPMJCR20T1 (Japan) and JST FOREST Grant No. JPMJFR2238 (Japan). This work was supported by Japan Science and Technology Agency (JST) as part of Adopting Sustainable Partnerships for Innovative Research Ecosystem (ASPIRE), Grant Number JPMJAP2426. Work at Princeton was supported by NSF, grant OAC-2118310, and by NSF through the Princeton Center for Complex Materials, a Materials Research Science and Engineering Center DMR-2011750.
@@
S I Density functional calculation
First-principles calculations are carried out using the WIEN2k software package, which employs the full-potential linearized augmented plane wave (FP-LAPW) method with local orbitals Blaha et al. (2020). The exchange-correlation energy is treated within the Generalized Gradient Approximation (GGA) using the functional formulated by Perdew, Burke, and Ernzerhof (PBE) Perdew et al. (1996) in the $P6/mmm$ structure. We employ a uniform grid of 4096 $k$-points to ensure accurate sampling of the Brillouin zone and relax the lattice constants $a$ and $c$. This yields $a=5.8260\mathrm{\AA},\ c=3.5666\mathrm{\AA}$ for LaRu3Si2 and $a=5.7967\mathrm{\AA},\ c=3.7538\mathrm{\AA}$ for hypothetical LaRu3Ge2. We calculate the band structures and the density of states in the presence of spin-orbit coupling. For the density of states (DOS) and partial density of states (PDOS), we divide the calculated values by $2$ to express them in units of $\mathrm{eV}^{-1}$ per unit cell (u.c.) per spin.
Fig. S1 shows our PDOS for the Ru-$4d_{z^{2}}$, Ru-$4d_{xy}$, and Ru-$4d_{xz}$ states of LaRu3Si2 and LaRu3Ge2. The symmetrically equivalent Ru-$4d_{x^{2}-y^{2}}$ and Ru-$4d_{yz}$ states are not shown separately. We provide projected band structures in the full Brillouin zone for LaRu3Si2 and LaRu3Ge2 in Fig. S2.
Figure S1: (color online). Partial density of states (PDOS) for Ru-$4d_{z^{2}}$, Ru-$4d_{xy}$, and Ru-$4d_{xz}$ states of LaRu3Si2 and LaRu3Ge2.
@@
S II Crystal growth
Crystals of LaRu3(Si1-xGex)2 (LRSG) are prepared by arc melting in argon atmosphere after careful evacuation to $10^{-3}\,$a. The ingot is turned and remelted at least three times, and weight loss is typically around $0.5\,\%$. To stabilize the LaRu3Si2 structure over the competing phase LaRu2Si2, excess Ru metal is added to the starting composition Barz (1980); Vandenberg and Barz (1980); Kishimoto et al. (2002); Li et al. (2011); Mielke et al. (2021). The solution limit of Ge is found to be $x=0.07$, and the required amount of Ru increases when approaching this value: We use a melt of starting composition LaRu4.5(Si1-xGex)2 for $x=0,0.02,0.05,$ and LaRu5(Si1-xGex)2 for $x=0.07$.
S III Powder X-ray diffraction
We perform powder X-ray diffraction (XRD) to confirm the formation of LaRu3(Si1-xGex)2 as well as the absence of the competing LaRu2Si2 (122). The data include reflections from nonmagnetic, elemental Ru, because Ru was added to the melt to destabilize the chemically stable 122 phase. In the main text, we neglect minor effects of subtle orthorhombic distortions in this compound. This approximation is justified by the Rietveld analysis, as the major reflections are well accounted for by space group $P6/mmm$.
@@
We perform magnetization measurements to detect the superconducting Meissner effect in a low magnetic field of $\mu_{0}H=0.05\,$T. Samples of mass $\sim 100\,$mg were mounted in a Quantum Design Magnetic Properties Measurement System (MPMS). We correct for the demagnetization effect in $\chi$ by approximating our roughly shaped crystals as a sphere, which has a demagnetization factor $N=1/3$. The intrinsic susceptibility is then obtained using the relation: $\chi=\chi_{\mathrm{exp}}/(1-N\chi_{\mathrm{exp}})$. Slightly smaller values than the ideal $\chi=-1$ (perfect Meissner shielding) may arise due to deviations from the spherical approximation.

### v2

# Chemical enhancement of superconductivity in LaRu3Si2 with mode-selective coupling between kagome phonons and flat bands

@@ Abstract
In kagome metals, flat electronic bands induced by frustrated hopping are a platform for strong electron correlations. In particular, a selective coupling of flat band states to certain kagome phonon modes is proposed as a universal origin of superconductivity in this material class. Here, we investigate the superconductivity in the kagome system LaRu3(Si1-xGex)2 by chemical pressure tuning while preserving the Ru-$4d$ states that constitute the kagome flat bands. We observe a sizable enhancement in the density of states up to $x=0.07$, as determined by the specific heat, with a concomitant increase in the superconducting transition temperature $T_{\mathrm{c}}$. Ge-dopants induce a uniaxial lattice expansion along the $c$-axis. Our first-principles calculations suggest that this mitigates the detrimental effect of hybridization between kagome layers and reduces the dispersion of the Ru-$d_{x^{2}-y^{2}}$ flat band. The calculated chemical potential moves closer to the maximum in the energy-dependent density of states. Our result is consistent with a theoretical prediction of tunable flat band superconductivity in LaRu3Si2 by mode-selective coupling between specific kagome phonons and the Ru-$d_{x^{2}-y^{2}}$ orbitals.
@@
The kagome lattice is a two-dimensional network of corner-sharing triangles with distinct properties in terms of its electronic band topology Yin et al. (2022); Wang et al. (2023); Checkelsky et al. (2024), electron-electron correlations, and phonon spectra Deng et al. (2025). This Letter discusses phonon-driven superconductivity in the intermetallic LaRu3Si2, where the kagome lattice is formed by Ru atoms as depicted in Fig. 1(a). Considering first a single orbital model on an isolated kagome lattice in Fig. 1(b), one obtains three bands forming a flat band and Dirac cone [Fig. 1(c)]. In particular, the dispersionless upper band originates from frustrated hopping between the three sites of the kagome lattice Bergman et al. (2008). Kane-Mele type spin-orbit coupling Kane and Mele (2005) further opens a gap and imparts a Chern number onto the flat band Guo and Franz (2009a); Bolens and Nagaosa (2019); Kang et al. (2020).
@@
The single-orbital model is not sufficient to describe the topological indices and band dispersion of real-world kagome metals Jiang et al. (2025), but such materials retain key features of this simplified electronic structure if the coupling between consecutive kagome layers is weak  Lin et al. (2018); Kang et al. (2019, 2020); Li et al. (2021); Regnault et al. (2022); Ye et al. (2024). As an example, Fig. 1(d) presents the electronic band structure of LaRu3Si2, which hosts an hourglass-shaped dispersion originating from Ru-$4d$ orbitals along the $\Gamma$$K$$M$ path. LaRu3Si2 is a rare kagome superconductor with a flat band near the Fermi level, yet without any magnetic instability of said flat band Mielke et al. (2021); Deng et al. (2025). In ab-initio theory, the flat band states couple selectively to certain acoustic kagome phonons Deng et al. (2025), discussed further at the end of this Letter, and realize the highest observed transition temperature $T_{\mathrm{c}}=6.7\,$K among the family of $RT_{3}M_{2}$ ($R$ = rare-earth, $T$ = transition metal and $M$ = main group element) Barz (1980); Ku et al. (1980); Chaudhary et al. (2023); Gui and Cava (2022); Rauchschwalbe et al. (1984); Mielke et al. (2021); Gong et al. (2022) and among kagome metals as a material class Deng et al. (2025). At higher temperature, other (optical) phonon modes of LaRu3Si2 soften and cause charge-density-wave (CDW) orders Deng et al. (2025), possibly accompanied by breaking of time-reversal symmetry Plokhikh et al. (2024); Mielke et al. (2024). The phenomenology is reminiscent of the kagome superconductors $AM_{3}$b5 ($A=$K, Cs, and Rb; $M=$V, Cr, …) Wilson and Ortiz (2024); Ortiz et al. (2019). In that material family, the most closely related compound is CsCr3Sb5, which hosts flat bands at $E_{\mathrm{F}}$ and a charge density wave (CDW) with the same modulation vector as LaRu3Si2 Liu et al. (2024a); however, all these phenomena occur at much lower temperature in CsCr3Sb5.
@@
Real kagome metals are three-dimensional, and their flat bands often exhibit strong dispersion along the $k_{z}$ direction of reciprocal space Kang et al. (2020); Huang et al. (2022); Ye et al. (2024). This behavior is also observed in LaRu3Si2 along the $\Gamma-A$ line in Fig. 1(d). Therefore, flat-band electrons are not fully localized in known bulk kagome metals. To achieve tailored flat bands, it has been proposed to either suppress interlayer coupling [$t_{\mathrm{out}}$ in Fig. 1(b)] by isolating kagome layers Jovanovic and Schoop (2022), or to introduce 3D frustrated hopping Wakefield et al. (2023). For the latter approach, 3D flat bands have been predicted for the pyrochlore lattice Guo and Franz (2009b), a 3D analogue of the kagome lattice, and have recently been observed Wakefield et al. (2023); Huang et al. (2024). Despite its simplicity, the former approach has not yet been conclusively implemented to enhance electron correlation phenomena and superconductivity, possibly driven by specific phonon modes, on the kagome lattice.
@@
Here, we engineer the electronic structure of LaRu3Si2 by substituting Si with the larger volume element Ge. We observe a uniaxial lattice expansion along the $c$-axis and a marked increase of the superconducting critical temperature $T_{\mathrm{c}}$. An enhancement of the total density of states (DOS) with Ge-doping underpins the observed evolution of $T_{\mathrm{c}}$, as confirmed by specific heat measurements. Our electronic structure calculations reproduce these behaviors; they suggest the bandwidth of the flat band decreases due to suppression of interlayer hopping $t_{\mathrm{out}}$ by the anisotropic lattice expansion. The calculations also suggest an effective hole doping through chemical pressure, moving the flat band closer to the Fermi level. The increase in $T_{\mathrm{c}}$ is broadly consistent with the scenario of mode-selective coupling to kagome phonons in LaRu3Si2 Deng et al. (2025). As summarized in Fig. 1(e), our results are in stark contrast to previous attempts to tune superconductivity in LaRu3Si2 by chemical substitution at the Ru site, which generally results in a sharp decrease of $T_{\mathrm{c}}$ Li et al. (2012, 2016); Chakrabortty et al. (2023).
@@
Crystals are obtained by the melting technique SI . In Fig. 2(a), the LaRu3Si2 structure is confirmed by powder X-ray diffraction (XRD). The absence of the LaRu2Si2 phase is confirmed. Figure 2(c,d) shows the hexagonal lattice constants of LaRu3(Si1-xGex)2 (LRSG) extracted from XRD and their evolution with $x$ based on the LeBail analysis in the $P6/mmm$ structure. This neglects minor effects of symmetry-lowering charge order in this family of compounds Plokhikh et al. (2024). We find a robust value of $a$ and a linear expansion of the out-of-plane lattice constant $c$, which in combination gives a linear volume expansion in Fig. 2(e). The latter is consistent with Vegard’s law, and in good agreement with the trend derived from structure relaxation calculations of LaRu3Si2 and LaRu3Ge2.
@@
Magnetization and specific heat measurements are performed in a commercial cryostat SI . To confirm the bulk superconducting state in LRSG, we show the magnetic susceptibility $\chi$ in Fig. 3(a). Above the transition temperature $T_{\mathrm{c}}$, $\chi$ is small and positive, consistent with the weak paramagnetic Pauli susceptibility of LaRu3Si2 Kishimoto et al. (2002). Strong diamagnetism accompanying the superconducting transition appears as a sharp drop in $\chi$ around $6\sim 7\,$K.
@@
To precisely determine $T_{\mathrm{c}}$, we measure the specific heat $C(T)$ in zero field and at 555\,5T, which is high enough to entirely suppress the superconductivity [Fig. 3(b)]. In the former traces, a sharp kink appears at $T_{\mathrm{c}}$, the second-order transition into the superconducting state. We model the normal state contribution $C_{\mathrm{N}}$ by electronic and phonon terms, viz. $C_{\mathrm{N}}(T)=\gamma T+\beta T^{3}+\eta T^{5}$ [lines in Fig. 3(b)]. The higher order phonon contribution $\eta T^{5}$ is necessary to describe the data  Li et al. (2011). After subtraction of $C_{\mathrm{N}}(T)$, we fit the superconducting specific heat $C(T)-C_{\mathrm{N}}(T)$ to the prediction of Bardeen-Cooper-Schrieffer (BCS) theory by equal-area construction in Fig. 3(c) Tinkham (2004), confirming an enhancement in $T_{\mathrm{c}}$ with $x$ in Fig. 3(d), reaching up to 7.17.17.1\,7.1K for $x=0.07$.
In Fig. 3(e), a positive slope of $\gamma(x)$ is reported. Error analysis and the evolution of $\beta$, from which we calculate $\Theta_{\mathrm{D}}$, are discussed in the Supplementary Information SI ; see also Fig. 4(f). Given the formula $\gamma=\pi^{2}k_{\mathrm{B}}^{2}N(E_{\mathrm{F}})/3$ with $N(E_{\mathrm{F}})$ being the density of states at the Fermi energy $E_{\mathrm{F}}$, the enhancement of $\gamma$ corresponds to an increase of $N(E_{\mathrm{F}})$ [Fig. 3(e)]. The trend is reproduced by our density functional theory (DFT) calculations, where $N(E_{\mathrm{F}})$ increases from $2.9\,\mathrm{/eV/u.c/spin}$ to $3.3\,\mathrm{/eV/u.c/spin}$ when going from LaRu3Si2 to LaRu3Ge2, c.f. Fig. 3(f). Calculations at an intermediate $x$ value are discussed in Ref. SI . We now take a closer look at the electronic bands to understand the behavior of $N(E_{\mathrm{F}},x)$. The projected band structures for LaRu3Ge2 and LaRu3Si2 are shown in Fig. 4(a),(b). Focusing on the $\Gamma-A$ line, which is parallel to the $k_{z}$-axis, we find that substitution of Si with Ge decreases the bandwidth of the flat band at $E_{\mathrm{F}}$; the uniaxial lattice expansion along the $c$-axis by Ge doping suppresses interlayer coupling between kagome layers.
In the partial density of states (PDOS) of LaRu3Ge2, Fig. 4(c),(d), the contribution of $4d_{z^{2}}$ grows weaker while the PDOS peak of $4d_{x^{2}-y^{2}}$ sharpens, corresponding to a narrowing of the flat band through chemical pressure tuning. Moreover, the calculations indicate a slight shift of the PDOS peak towards $E_{\mathrm{F}}$. From the viewpoint of superconductivity, Ge doping has the favorable effect of reducing the bandwidth, but it also helps to drive $E_{\mathrm{F}}$ into the fractionally filled flat band. Note that recent theory work predicts an enhancement of $T_{\mathrm{c}}$ by hole doping LaRu3Si2 Deng et al. (2025).
To discuss these results, we consider the McMillan formalism for BCS,
$T_{\mathrm{c}}=\frac{\left<\omega\right>}{1.20}\exp\left(\frac{-1.04(1+\lambda)}{\lambda-\mu^{*}(1+0.62\lambda)}\right),$ (1)
where $\omega\right>$ is a characteristic phonon frequency after weighting by the electron-phonon interaction SI and $\mu^{*}$ is the Coulomb pseudopotential McMillan (1968). The electron-phonon coupling constant $\lambda$ can be expressed as
$\lambda=2\frac{N(E_{\mathrm{F}})}{N}\frac{\hbar\langle g^{2}\rangle}{\hbar^{2}\langle\omega^{2}\rangle},$ (2)
where $N$ and $\langle g^{2}\rangle$ represent the number of sites and the weighted phonon linewidth  McMillan (1968). For LaRu3Si2 we expect $\lambda<1$ based on (i) recent ab-initio modeling, which predicts $\lambda\approx 0.8$, and (ii) $\Delta C/(\gamma T_{\mathrm{c}})$ in Fig. 4(h), which is characteristic of moderate coupling strength.
According to recent theory, $d_{x^{2}-y^{2}}$ flat bands have a mode-selective electron-phonon coupling to the $B_{3u}$ kagome phonon Deng et al. (2025). This low-energy mode corresponds to a planar oscillation into and out of the large hexagonal plaquette in the kagome lattice, illustrated in Fig. 4(e). In the scenario of Deng et al. Deng et al. (2025), the widely used Dynes (1972)
$\omega\right>\approx 0.828\,\Theta_{\mathrm{D}}$ (3)
[…]
In summary, we control superconductivity in LaRu3Si2 by chemical pressure and use specific heat measurements – which is not accessible for recent studies of enhanced superconductivity in LaRu3Si2 under high external pressure Ma et al. (2024); Li et al. (2025) – to track the electronic density of states and phononic Debye temperature. Although chemical pressure tuning has been implemented in a number of superconductors SI , simultaneous reports of $T_{\mathrm{c}}$ and $\gamma$ are less common, and none such has targeted kagome metals. Tuning the flat band by uniaxial strain, using large single crystals, remains a challenge for future research.
@@
Acknowledgements We thank T.-h. Arima for enlightening discussions. This work was supported by JSPS KAKENHI Grants Nos. JP21K13877, JP22H04463, JP22K20348, JP23H05431, 23K13057, 24H01607, and 24H01604, as well as JST CREST Grant Nos. JPMJCR1874 and JPMJCR20T1 (Japan) and JST FOREST Grant No. JPMJFR2238 (Japan). This work was supported by Japan Science and Technology Agency (JST) as part of Adopting Sustainable Partnerships for Innovative Research Ecosystem (ASPIRE), Grant Number JPMJAP2426. R.Y. acknowledges support from the US National Science Foundation (NSF) Grant Number 2201516 under Accelnet program of Office of International Science and Engineering (OISE). Work at Princeton was supported by NSF, grant OAC-2118310, and by NSF through the Princeton Center for Complex Materials, a Materials Research Science and Engineering Center DMR-2011750.
@@
S I Crystal growth
Crystals of LaRu3(Si1-xGex)2 (LRSG) are prepared by arc melting in an argon atmosphere after careful evacuation to $10^{-3}\,$a. The ingot is turned and remelted at least three times, and weight loss is typically around $0.5\,\%$. To stabilize the LaRu3Si2 structure over the competing phase LaRu2Si2, excess Ru metal is added to the starting composition Barz (1980); Vandenberg and Barz (1980); Kishimoto et al. (2002); Li et al. (2011); Mielke et al. (2021). The solution limit of Ge is found to be $x=0.07$, and the required amount of Ru increases when approaching this value: We use a melt of starting composition LaRu4.5(Si1-xGex)2 for $x=0,0.02,0.05,$ and LaRu5(Si1-xGex)2 for $x=0.07$.
S II Powder X-ray diffraction
We perform powder X-ray diffraction (XRD) to confirm the formation of LRSG as well as the absence of the competing LaRu2Si2. The data include reflections from nonmagnetic, elemental Ru, because Ru is added to the melt to destabilize the chemically stable LaRu2Si2 phase.
S III Density functional calculation
First-principles calculations are carried out using the WIEN2k software package, which employs the full-potential linearized augmented plane wave (FP-LAPW) method with local orbitals Blaha et al. (2020). The exchange-correlation energy is treated within the Generalized Gradient Approximation (GGA) using the functional formulated by Perdew, Burke, and Ernzerhof (PBE) Perdew et al. (1996) in the $P6/mmm$ structure. We employ a uniform grid of 4096 $k$-points to ensure accurate sampling of the Brillouin zone and relax the lattice constants $a$ and $c$. This yields $a=5.8260\mathrm{\AA},\ c=3.5666\mathrm{\AA}$ for LaRu3Si2 and $a=5.7967\mathrm{\AA},\ c=3.7538\mathrm{\AA}$ for hypothetical LaRu3Ge2. We calculate the band structures and the density of states in the presence of spin-orbit coupling (SOC). For the density of states (DOS) and partial density of states (PDOS), we divide the calculated values by $2$ to express them in units of $\mathrm{eV}^{-1}$ per unit cell (u.c.) per spin. A previous report on the band structure of LaRu3Si2 assumes the $P6_{3}/m$ structure with two time larger unit cell Mielke et al. (2021). However, a recent study shows that the proper structure is orthorhombic Plokhikh et al. (2024). Due to the different structure and band folding in the $P6_{3}/m$ space group, slight differences can be seen in our calculation results compared to the previous study. On the other hand, our calculations are largely consistent with recent theoretical work on LaRu3Si2 Deng et al. (2025). For our calculations, we ignore the structural transition from $P6/mmm$ to $Cccm$ and subsequent CDW transitions Plokhikh et al. (2024). Recent theoretical work shows that the flat band persists in the orthorhombic $Cccm$ phase Deng et al. (2025).
Fig. 4 shows the PDOS for the Ru $4d$ orbitals of LaRu3Si2 and LaRu3Ge2. We provide projected band structures in the full Brillouin zone for LaRu3Si2 and LaRu3Ge2 in Fig. S1.
To check the impact of spin-orbit coupling (SOC), we compare band structure and DOS with and without SOC in Fig. S2. For both cases, the flat band is present at the Fermi energy, and LaRu3Ge2 has a larger DOS than LaRu3Si2.
In addition to LaRu3Si2 and LaRu3Ge2, we illustrate in Fig. S3 the calculations performed for LRSG with $x=8\,\%$ by the virtual crystal approximation (VCA) Bellaiche and Vanderbilt (2000) implemented in VASP Kresse and Furthmüller (1996). The DOS systematically increases around the Fermi level upon doping Ge [Fig. S3(a)].
@@
Figure S2: (color online). Comparison of ab initio calculations with and without SOC. (a),(b) Band structure for LaRu3Si2 with and without spin-orbit coupling (SOC). (c),(d) Band structure for LaRu3Ge2 with and without SOC. (e) Density of states (DOS) for LaRu3Si2 with and without SOC. (f) DOS for LaRu3Ge2 with and without SOC. (g) Comparison of LaRu3Si2 and LaRu3Ge2 with SOC. (h) Comparison of LaRu3Si2 and LaRu3Ge2 without SOC.
[…]
@@
We perform magnetization measurements to detect the superconducting Meissner effect in a low magnetic field of $\mu_{0}H=0.05\,$T. Samples of mass $\sim 100\,$mg are mounted in a Quantum Design Magnetic Properties Measurement System (MPMS). We correct for the demagnetization effect in $\chi$ by approximating our roughly shaped crystals as a sphere, which has a demagnetization factor $N=1/3$. The intrinsic susceptibility is then obtained using the relation: $\chi=\chi_{\mathrm{exp}}/(1-N\chi_{\mathrm{exp}})$. Slightly smaller values than the ideal $\chi=-1$ (perfect Meissner shielding) may arise due to deviations from the spherical approximation.